Write an equation for the parabola whose vertex is at (-8,4) and passes through (-6,2).
Answer by venugopalramana(3286)

Write an equation for the parabola whose vertex is at (-8,4) and passes through (-6,2).
EQN.OF PARABOLA IN STD FORM IS
(Y-K)^2=4A(X-H)
WHERE (H,K) IS VERTEX...H=-8...K=4
EQN IS
(Y-4)^2=4A(X+8)...IT PASSES THROUGH (-6,2)..SO,...
(2-4)^2=4A(-6+8)=8A
4=8A
A=0.5
EQN.IS
(Y-4)^2=2(X+8)